A daredevil decides to jump a canyon. It's walls are equally high and 12 m apart. He takes off by driving a motorcycle up a short ramp sloped at an angle of 19°. What minimum speed must he have in order to clear the canyon?
1
Expert's answer
2019-10-10T09:53:49-0400

Let us use notation "\theta = 19^{\circ}", "L = 12 m".

The horizontal distance, covered by the projectile aimed at an angle "\theta" with initial speed "v_0" is "S = \frac{v_0^2 \sin 2 \theta}{g}". This distance must be minimum "L" in order for a daredevil not to fall. Hence, from the previous equation, "\frac{v_0^2 \sin 2 \theta}{g} = L \Rightarrow v_0 = \sqrt{\frac{ g L}{ \sin 2 \theta}} \approx 13.83 \frac{m}{s}".